# Augmenting Human Expertise in Weighted Ensemble Simulations through Deep Learning based Information Bottleneck

### v1

@@ II.5 Simulation Details
Figure 4: Average free energy surfaces of alanine dipeptide from three separate 10 $\mu s$ runs of (b,e) the WE solely based on expert-based CVs and (c,f) the hybrid SPIB-WE, compared with the reference 4 $\mu s$ brute-force simulations (a,d).
@@
To demonstrate our hybrid SPIB-WE approach, we benchmark it using two model systems: the small biomolecule alanine dipeptide in water and the mini folding chignolin mutant protein, CLN025, in implicit solvent. All simulations are performed with the OpenMM simulation packageEastman et al. (2017) patched with PLUMED 2.7Tribello et al. (2014) through the OpenMM-PLUMED interface plugin.Eastman (2021)
@@
For alanine dipeptide, the PDB file was obtained from http://ftp.imp.fu-berlin.de/pub/cmb-data/alanine-dipeptide-nowater.pdb. The system was prepared using OpenMM with the AMBER ff14SB force field,Maier et al. (2015) and subjected to explicit solvation using the TIP3P water model. The MD integrations are performed with a timestep of 2 fs. The system is first equilibrated under the constant number, volume, and temperature (NVT) ensemble for 10 ns, followed by the equilibration in the constant number, pressure, and temperature (NPT) ensemble for 10 ns. The resulting structure after this equilibration phase serves as the starting structure for all subsequent WE simulations. The NPT ensemble is finally employed for the WE simulations. We maintain a constant temperature of 300 K using the Langevin integrator with a collision frequency of 1 ps-1, while the pressure is maintained at 1 bar using the Monte Carlo barostat.Åqvist et al. (2004)
@@
We obtained the folded structure of CLN025 (amino acid sequence YYDPETGTWY) from the 2RVD entry of the Protein Data Bank. The system was prepared using OpenMM with the AMBER ff14SB-onlysc force field,Maier et al. (2015) and subjected to generalized Born (GB) implicit solvation using the GBn2 model.Nguyen, Roe, and Simmerling (2013) The MD integrations are performed with a timestep of 2 fs. The system is equilibrated under the NVT ensemble for 10 ns. The simulations are performed using the Langevin integrator at the melting temperature of 340 K with a collision frequency of 1 ps-1. The resulting structure after equilibration is utilized as the starting structure for all WE simulations. The WE simulations for CLN025 are performed in the NVT ensemble.
@@
All WE simulations are carried out using the open-source WESTPA 2.0 software package.Russo et al. (2022) All simulation trajectories are saved with a time resolution of 1 ps for both systems, facilitating SPIB training, as well as rate and free energy profile estimation. The resampling time $\tau$ is set to 20 ps and the target number of replicas per bin is set to 4, identical for all WE and SPIB-WE simulations. We initially perform $N_{0}=20$ WE iterations for alanine dipeptide, and $N_{0}=40$ WE iterations for CLN025 before starting SPIB training. We use same SPIB update interval ($N_{\text{update}}=20$ WE iterations) for both systems. We observe that SPIB training time remains nearly unchanged as the training dataset size increases. Therefore, we train SPIB on data from the last $N_{\text{last}}=500$ WE iterations for both systems. We set the lag time of SPIB as the half of the WE resampling time, resulting in $\Delta t=10$ ps. For binning the SPIB-learned CVs, we set the desired number of occupied SPIB bins to 100 for both systems, ensuring a relatively fine grid along the 2D CVs while balancing computational costs.
@@
We find that due to the relatively slow transition rates, some explored states in the initial stages of the SPIB-WE process may have very small weights. Considering these weights during SPIB training at the outset may lead SPIB to overlook these states. Typically, SPIB tends to disregard states with populations smaller than $10^{-3}$. Therefore, in this work, we ignore the weights during the first 100 WE iterations. After this initial period, all weights are considered as usual. As a general guideline for future users, it’s recommended to ignore weights in the initial stages of the SPIB-WE protocol. If SPIB identifies too many states or if states of interest already have significant weights ($>10^{-3}$), then weights are recommended to be considered. This allows SPIB to ignore states with very small weights and focus on sampling transitions between more significant states.
@@
Table 1: Rate Constants (in ns-1) between Different Metastable States for alanine dipeptide. aIn the case of the brute-force simulation, the first value indicates the average rate constant, while the values within the brackets indicate the 95% confidence interval computed using Bayesian bootstrapping. bFor both the SPIB-WE and expert-based WE methods, the reported values represent the mean along with the standard error of the mean derived from three independent runs.
@@
Figure 6: Evolution of rate constants over aggregate simulation time for the expert-based WE simulations (red), and the SPIB-WE approach (blue). The error bars represent the standard error of the mean from three independent runs. The black line shows the ground-truth obtained from the 4 $\mu$s brute-force simulation, and the shaded area shows the 95% confidence interval estimated using Bayesian bootstrapping.
@@ III.1 Alanine Dipeptide
We identify three metastable states in alanine dipeptide: $\alpha_{R}$, $\alpha_{L}$, and $P_{II}$, as defined in Fig. 3(b). Subsequently, we assess both the expert-based WE and the SPIB-WE by estimating rate constants between these states, as summarized in Table 1. Fig. 6 shows the evolution of the estimated rate constants over aggregate simulation time for the expert-based WE and SPIB-WE. Although the total simulation time to converge is similar to that of brute-force unbiased reference MD, it is important to note that WE runs simulations in parallel as a series of short trajectories. This parallelization significantly accelerates the convergence process, making WE more efficient despite the comparable total simulation time. In this case, since the expert-based CVs $\phi$ and $\psi$ effectively characterize the dynamics of alanine dipeptide in water, both the expert-based WE and the SPIB-WE accurately estimate the rate constants between the regions of interest. We note comparable convergence and accuracy of the kinetic rates for both methods, although the expert-based WE tends to exhibit smaller variance between different runs. These results serve as a validation that SPIB can learn CVs as effective as $\phi$ and $\psi$ from pairwise heavy-atom distances for the alanine dipeptide systems. For more complex systems, such as chignolin, where expert-based CVs struggle to capturing its dynamics, SPIB’s advantage in managing intricate dynamics becomes evident, as illustrated in Figure 10.
@@ III.2 Chignolin mutant CLN025
The chignolin mutant, CLN025, with a sequence of YYDPETGTWY is one of the smallest fast-folding proteinsHonda et al. (2008) that folds into a $\beta$-hairpin as shown in Fig. 8(a). For this system, the commonly expert-based CVs for WE simulations are mass-weighted RMSD to the folded structure (PDB ID:2RVD) and mass-weighted Rg.Ojha et al. (2023); Ahn et al. (2021) We place both CVs evenly in bins of 0.03 nm in the range of [0.04 nm, 0.70 nm] for RMSD and in the range of [0.41 nm, 0.83 nm] for Rg. For SPIB training, nearest-neighbor heavy-atom distances between all residues separated by two or more residues serve as the SPIB input, resulting in 28 distances in total. The WE solely based on these two expert-based CVs is executed for three independent runs of 10 $\mu$s each. Similarly, the hybrid SPIB-WE approach is conducted for three independent runs of approximately 10 $\mu$s each. Additionally, three brute-force reference unbiased 4 $\mu$s simulations were carried out to obtain reference values.
@@
Table 2: Rate Constants (in ns-1) between different metastable states for CLN025. aIn the case of the brute-force simulations, the first value indicates the average rate constant, while the values within the brackets indicate the 95% confidence interval computed using Bayesian bootstrapping. bFor both the SPIB-WE and expert-based WE methods, the reported values represent the mean along with the standard error of the mean derived from three independent runs.
@@
Since no bias is introduced in either the expert-based WE or SPIB-WE, we attribute SPIB-WE’s superior performance to its enhanced capability to distinguish between different metastable configurations and pinpoint crucial state-to-state transitions. This increased discriminatory power leads to a higher number of uncorrelated successful transition events, thereby enhancing SPIB-WE’s effectiveness over expert-based WE in capturing state-to-state transitions and reducing correlations arising from resampling processes.
Figure 11: Evolution of rate constants over aggregate simulation time for the WE simulations (red), and the SPIB-WE approach (blue). For the WE and SPIB-WE methods, the error bars represent the standard error of the mean from three independent runs. The black line shows the ground-truth obtained from the long brute-force simulations, and the shaded area shows the 95% confidence interval estimated from 3 independent runs. The first 1 $\mu$s simulation is excluded from the rate analysis.
@@
For a more in-depth analysis of the SPIB-learned CVs, we aggregate data from the last 500 WE iterations of three independent SPIB-WE runs and retrain a single SPIB model, as depicted in Fig. 12. Fig. 12(a) showcases the free energy surface (FES) of the learned CVs, while Fig. 12(b) illustrates the metastable states identified by the algorithm. Utilizing a lag time $\Delta t=10$ ps, we identify 13 metastable states from the data. The conformational ensembles and associated transitions for these 13 macrostates are visually represented in Fig. 12(c). S1010{10}10 and S1212{12}12 correspond to the folded state, while S22{2}2, S33{3}3, and S66{6}6 represent the collapsed unfolded states. We also observe a misfolded-like conformation in S55{5}5, where CLN025 adopts an $\alpha$-turn from Asp3 to Gly7 instead of the native $\pi$-turn from Asp3 to Thr8. However, as suggested by the FES in Fig. 9, this misfolded conformation is not very stable in implicit solvent simulations, consistent with findings from the previous study.Chen et al. (2021) Interestingly, our SPIB-WE protocol is also able to sample conformations where CLN025 forms a helix structure, as seen in S00{0}, S77{7}7, and S9. As all these identified states have significant Boltzmann weights, we indeed observe that they are also sampled by expert-based WE. However, since the FES of the expert-based CVs shows no distinct basins, it would be more challenging to sample and identify these states using expert-based CVs. This result showcases how SPIB automatically clusters sampled MD conformations into important metastable states and maps all key state-to-state transitions into a low-dimensional CV space. Therefore, by utilizing these learned CVs to guide sampling, the WE simulations can better focus on important state-to-state transitions, thus achieving improved performance.
@@ IV Discussion
In this work, we exclusively tested our method on well-studied model systems. However, we anticipate that using more expert-based CVs will be beneficial for tackling more challenging systems. In such cases, users can adjust the expert-based CVs whenever exploration within the current set is hindered. While this approach may require more manual intervention, it provides users with greater flexibility, leveraging all available expert-based CVs to guide exploration effectively. Moreover, we employ a rectilinear grid with equal bin sizes for binning both the expert-based and SPIB-learned CVs for simplicity, but more advanced binning schemes could further enhance sampling efficiency.Torrillo, Bogetti, and Chong (2021) Additionally, to accelerate the convergence of WE simulations, it is also promising to initialize the WE simulations with better weighted initial configurations, as demonstrated in previous studies.Ojha et al. (2023); Ahn et al. (2021) These initial configurations can be selected by training a SPIB on short unbiased or biased trajectories, similar to the approach used in Ref. Ojha et al., 2023. Finally, considering SPIB’s utility in constructing MSMs, another natural way to speed up the convergence of rate and FES estimation is through employing more efficient analysis tools like MSM and haMSM to analyze the WE data. We leave the exploration of these extension directions for future work.
@@
See supplementary material for other numerical details.

### v2

@@ II.5 Simulation Details
To demonstrate our hybrid SPIB-WE approach, we benchmark it using two model systems: the small biomolecule alanine dipeptide in water and the mini folding chignolin mutant protein, CLN025, in implicit solvent. All simulations are performed with the OpenMM simulation package[40] patched with PLUMED 2.7[41] through the OpenMM-PLUMED interface plugin.[42]
@@
For alanine dipeptide, the PDB file was obtained from http://ftp.imp.fu-berlin.de/pub/cmb-data/alanine-dipeptide-nowater.pdb. The system was prepared using OpenMM with the AMBER ff14SB force field,[43] and subjected to explicit solvation using the TIP3P water model. The MD integrations are performed with a timestep of 2 fs. The system is first equilibrated under the constant number, volume, and temperature (NVT) ensemble for 10 ns, followed by the equilibration in the constant number, pressure, and temperature (NPT) ensemble for 10 ns. The resulting structure after this equilibration phase serves as the starting structure for all subsequent WE simulations. The NPT ensemble is finally employed for the WE simulations. We maintain a constant temperature of 300 K using the Langevin integrator with a collision frequency of 1 ps-1, while the pressure is maintained at 1 bar using the Monte Carlo barostat.[44]
@@
We obtained the folded structure of CLN025 (amino acid sequence YYDPETGTWY) from the 2RVD entry of the Protein Data Bank. The system was prepared using OpenMM with the AMBER ff14SB-onlysc force field,[43] and subjected to generalized Born (GB) implicit solvation using the GBn2 model.[45] The MD integrations are performed with a timestep of 2 fs. The system is equilibrated under the NVT ensemble for 10 ns. The simulations are performed using the Langevin integrator at the melting temperature of 340 K with a collision frequency of 1 ps-1. The resulting structure after equilibration is utilized as the starting structure for all WE simulations. The WE simulations for CLN025 are performed in the NVT ensemble.
@@
All WE simulations are carried out using the open-source WESTPA 2.0 software package.[14] All simulation trajectories are saved with a time resolution of 1 ps for both systems, facilitating SPIB training, as well as rate and free energy profile estimation. The resampling time $\tau$ is set to 20 ps and the target number of replicas per bin is set to 4, identical for all WE and SPIB-WE simulations. We initially perform $N_{0}=20$ WE iterations for alanine dipeptide, and $N_{0}=40$ WE iterations for CLN025 before starting SPIB training. We use same SPIB update interval ($N_{\text{update}}=20$ WE iterations) for both systems. We observe that SPIB training time remains nearly unchanged as the training dataset size increases. Therefore, we train SPIB on data from the last $N_{\text{last}}=500$ WE iterations for both systems. We set the lag time of SPIB as the half of the WE resampling time, resulting in $\Delta t=10$ ps. For binning the SPIB-learned CVs, we set the desired number of occupied SPIB bins to 100 for both systems, ensuring a relatively fine grid along the 2D CVs while balancing computational costs.
@@
We find that due to the relatively slow transition rates, some explored states in the initial stages of the SPIB-WE process may have very small weights. Considering these weights during SPIB training at the outset may lead SPIB to overlook these states. Typically, SPIB tends to disregard states with populations smaller than $10^{-3}$. Therefore, in this work, we ignore the weights during the first 100 WE iterations. After this initial period, all weights are considered as usual. As a general guideline for future users, it is recommended to ignore weights in the initial stages of the SPIB-WE protocol. If SPIB identifies too many states or if states of interest already have significant weights ($>10^{-3}$), then weights are recommended to be considered. This allows SPIB to ignore states with very small weights and focus on sampling transitions between more significant states.
@@
Table 1: Rate Constants (in ns-1) between Different Metastable States for alanine dipeptide. aIn the case of the brute-force simulation, the first value indicates the average rate constant, while the values within the brackets indicate the 95% credible interval computed using Bayesian bootstrapping. bFor both the SPIB-WE and expert-based WE methods, the reported values represent the mean along with the standard error of the mean derived from three independent runs.
@@
Figure 7: Evolution of rate constants over aggregate simulation time for the expert-based WE simulations (red), and the SPIB-WE approach (blue). The error bars represent the standard error of the mean from three independent runs. The black line shows the ground-truth obtained from the 4 $\mu$s brute-force simulation, and the shaded area shows the 95% credible interval estimated using Bayesian bootstrapping.
@@ III.1 Alanine Dipeptide
We identify three metastable states in alanine dipeptide: $\alpha_{R}$, $\alpha_{L}$, and $P_{II}$, as defined in Fig. 4(b). Subsequently, we assess both the expert-based WE and the SPIB-WE by estimating rate constants between these states, as summarized in Table 1. Fig. 7 shows the evolution of the estimated rate constants over aggregate simulation time for the expert-based WE and SPIB-WE. Although the total simulation time to converge is similar to that of brute-force unbiased reference MD, it is important to note that WE runs simulations in parallel as a series of short trajectories. This parallelization with resampling significantly accelerates the convergence process, making WE more efficient despite the comparable total simulation time. In this case, since the expert-based CVs $\phi$ and $\psi$ effectively characterize the dynamics of alanine dipeptide in water, both the expert-based WE and the SPIB-WE accurately estimate the rate constants between the regions of interest. We note comparable convergence and accuracy of the kinetic rates for both methods, although the expert-based WE tends to exhibit smaller variance between different runs. These results serve as a validation that SPIB can learn CVs as effective as $\phi$ and $\psi$ from pairwise heavy-atom distances for the alanine dipeptide systems. For more complex systems, such as chignolin, where expert-based CVs struggle to capturing its dynamics, SPIB’s advantage in managing intricate dynamics becomes evident, as illustrated in Figure 10.
@@ III.2 Chignolin mutant CLN025
The chignolin mutant, CLN025, with a sequence of YYDPETGTWY is one of the smallest fast-folding proteins[46] that folds into a $\beta$-hairpin as shown in Fig. 9(a). For this system, the commonly expert-based CVs for WE simulations are mass-weighted RMSD to the folded structure (PDB ID:2RVD) and mass-weighted Rg.[21, 20] We place both CVs evenly in bins of 0.03 nm in the range of [0.04 nm, 0.70 nm] for RMSD and in the range of [0.41 nm, 0.83 nm] for Rg. For SPIB training, nearest-neighbor heavy-atom distances between all residues separated by two or more residues serve as the SPIB input, resulting in 28 distances in total. The WE solely based on these two expert-based CVs is executed for three independent runs of 10 $\mu$s each. Similarly, the hybrid SPIB-WE approach is conducted for three independent runs of approximately 10 $\mu$s each. Additionally, three brute-force reference unbiased 4 $\mu$s simulations were carried out to obtain reference values.
@@
Table 2: Rate Constants (in ns-1) between different metastable states for CLN025. aIn the case of the brute-force simulations, the first value indicates the average rate constant, while the values within the brackets indicate the 95% credible interval computed using Bayesian bootstrapping. bFor both the SPIB-WE and expert-based WE methods, the reported values represent the mean along with the standard error of the mean derived from three independent runs.
@@
Since no bias is introduced in either the expert-based WE or SPIB-WE, we attribute SPIB-WE’s superior performance to its enhanced capability to distinguish between different metastable configurations and pinpoint crucial state-to-state transitions. This increased discriminatory power leads to a higher number of uncorrelated successful transition events, thereby enhancing SPIB-WE’s effectiveness over expert-based WE in capturing state-to-state transitions and reducing correlations arising from resampling processes. This, in turn, may explain the reduction in run-to-run variances of rate esimation and the faster convergence of the FES and rates from SPIB-WE.
Figure 12: Evolution of rate constants over aggregate simulation time for the WE simulations (red), and the SPIB-WE approach (blue). For the WE and SPIB-WE methods, the error bars represent the standard error of the mean from three independent runs. The black line shows the ground-truth obtained from the long brute-force simulations, and the shaded area shows the 95% credible interval estimated from 3 independent runs. The first 1 $\mu$s simulation is excluded from the rate analysis.
@@
For a more in-depth analysis of the SPIB-learned CVs, we aggregate data from the last 500 WE iterations of three independent SPIB-WE runs and retrain a single SPIB model, as depicted in Fig. 13. Fig. 13(a) showcases the free energy surface (FES) of the learned CVs, while Fig. 13(b) illustrates the metastable states identified by the algorithm. Utilizing a lag time $\Delta t=10$ ps, we identify 13 metastable states from the data. The conformational ensembles and associated transitions for these 13 macrostates are visually represented in Fig. 13(c). S1010{10}10 and S1212{12}12 correspond to the folded state, while S22{2}2, S33{3}3, and S66{6}6 represent the collapsed unfolded states. We also observe a misfolded-like conformation in S55{5}5, where CLN025 adopts an $\alpha$-turn from Asp3 to Gly7 instead of the native $\pi$-turn from Asp3 to Thr8. However, as suggested by the FES in Fig. 10, this misfolded conformation is not very stable in implicit solvent simulations, consistent with findings from the previous study.[47] Interestingly, our SPIB-WE protocol is also able to sample conformations where CLN025 forms a helix structure, as seen in S0, S7, and S9. As all these identified states have significant Boltzmann weights, we indeed observe that they are also sampled by expert-based WE. However, since the FES of the expert-based CVs shows no distinct basins, it would be more challenging to sample and identify these states using expert-based CVs. This result showcases how SPIB automatically clusters sampled MD conformations into important metastable states and maps all key state-to-state transitions into a low-dimensional CV space. Therefore, by utilizing these learned CVs to guide sampling, the WE simulations can better focus on important state-to-state transitions, thus achieving improved performance.
@@ IV Discussion
In this work, we exclusively tested our method on well-studied model systems. However, we anticipate that using more expert-based CVs will be beneficial for tackling more challenging systems. In such cases, users can adjust the expert-based CVs whenever exploration within the current set is hindered. While this approach may require more manual intervention, it provides users with greater flexibility, leveraging all available expert-based CVs to guide exploration effectively. Moreover, we employ a rectilinear grid with equal bin sizes for binning both the expert-based and SPIB-learned CVs for simplicity, but more advanced binning schemes could further enhance sampling efficiency.[18] Additionally, to accelerate the convergence of WE simulations, it is also promising to initialize the WE simulations with better weighted initial configurations, as demonstrated in previous studies.[21, 20] These initial configurations can be selected by training a SPIB on short unbiased or biased trajectories, similar to the approach used in Ref. 21. Finally, considering SPIB’s utility in constructing MSMs,[24] another natural way to speed up the convergence of rate and FES estimation is through employing more efficient analysis tools like MSM and haMSM to analyze the WE data. We leave the exploration of these extension directions for future work.
@@
See supplementary material for further numerical details.